OpenStudy (anonymous):

find the sum of the finite geometric series by using the formula for Sn 1-1/3+1/9-1/27+1/81-1/243+1/729

OpenStudy (anonymous):

Do you mean infinite? otherwise you could just add them together.

OpenStudy (anonymous):

infinite

OpenStudy (anonymous):

finite geometric series

OpenStudy (anonymous):

The equation to solve a geometric progression: \[a(1-r^{n+1})/(1-r) = \sum_{n}^{0}ar^n\] Since a = 1, r = -1/3, and n = 6: \[1(1-(-1/3)^7)/(4/3) = 0.750342936\]